OpenStudy (anonymous):

Find the length of the radius for this circle. (x - 3)² + (y - 5)² = 8

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=radius^2

OpenStudy (anonymous):

The standard form of a circle is: \[(x-h)^2 + (y-k)^2 = r^2\] So in this case, \(r^2\) = 8... What is r?

OpenStudy (anonymous):

That's the general formula for a circle

OpenStudy (anonymous):

2rad2 would be the legnth of the radius